OpenStudy (anonymous):

In how many ways can a teacher arrange 5 students in the front row of a classroom with a total of 40 students?

OpenStudy (anonymous):

Just make 5 slots.Put 20 in first slot,19 in second slot,18 in 3rd slot,17 in 4th slot,16 in 5th slot.Just find out the product of 20.19.18.17.16.that is answer.

OpenStudy (anonymous):

i dont think thats it, that gives me 1860480

OpenStudy (anonymous):

answer choices are 68,960,960 78,960,960 88,960,960 98,960,960

OpenStudy (anonymous):

Since order matters this ia a permutation 40P5 I am going to finish the problem (my way) and then show you my fast method. Pardon me if you already know it. 40P5=(40)(39)(38)(37)(36)=78960960

OpenStudy (anonymous):

n= total number of objects you are choosing from r= the number of objects you are choosing ................n! nCr=************* ...........(n-r)!*r! BUT I dont use it........to slow Start with how I do a Permutation 10P2=..................start with 10 and count down 2 numbers 10P2=10*9...............2 numbers because of the 2 after the P 10P2=10*9=90 8P3=8*7*6=336 Now the Combination starts off the SAME way 10C2=10*9 EXCEPT now you have a denominator ......................to get the denominator you start with 1 and count up to 2 ......................................... to whatever the number is following the C .............10*9 10C2=*******=...............this how you set it up. Now either cancel or multiply ..............1*2 ............10*9 10C2=******=90/2 =45 .............1*2 ..........8*7*6 8C3=*******=.....cancel the 2*3 in the den. with the 6 in the num,.....=8*7=56 ..........1*2*3 Special Note::::Only on combinations (not permutations) may you do the following .............10*9*8*7*6*5*4 10C7=*******************.....NO, do not continue .............1*2*3*4*5*6*7 Subtract......... 10 minus 7..........answer is 3 Replace the 7 with a 3 ............10*9*8 10C3=********=10*3*4..(i reduced/canceled)..=120 .............1*2*3